OpenStudy (anonymous):

Tickets to a concert cost $18 in advance or $25 on the day of the performance. A total of 940 tickets are sold, and the total receipts were $21,820. How many of each kind of ticket were sold? I need to try and find the two equations describing this relationship. I am stuck...

OpenStudy (theyankee):

Do you know how to solve equations with two variables?

OpenStudy (anonymous):

As in, having an "x" and a "y" within the problem? Yes. :) I'm just having a hard time taking the information given, and putting it together in a solvable problem. Once I get to that point, I think I am good.

OpenStudy (theyankee):

Okay ^^ Here's the equation to solve it, then 18x+25y=21,820 Where x is the amount tickets sold in advanced and y is the number of tickets sold at the gate. Hope it helps :)

OpenStudy (anonymous):

Thank you! I had thought that was the equation, but I wasn't 100% sure. It seemed too simple at first haha. :)

OpenStudy (anonymous):

Not to be a bother, but I am told that there should be two equations to get from this. At some point, do I need to be having a separate equation for the total number of tickets sold?

OpenStudy (theyankee):

Yes. You'll need to solve for x, then substitute the x value into the equation and resolve for y :)

OpenStudy (anonymous):

When I solve for x, I am getting -25y/18 + 10910/9... I feel like this isn't right...

OpenStudy (theyankee):

I don't know why you'd get something incorrect :/ I've been doing this for a while. Give me a minute and I'll look into it.

OpenStudy (anonymous):

Not a problem. I am quickly finding out that math is still not my forte almost 6 years later haha!

OpenStudy (theyankee):

Does your question happen to give the total number of tickets sold? Like, maybe 940?

OpenStudy (theyankee):

I consulted friends, ha, ha... Here's what I have for you. 240 sold in advance 700 sold day of concert

OpenStudy (anonymous):

Haha! Thank you! Does your friends happen to know the two equations that were used to solve this problem? I have a feeling that the first one is right, but I think there is a second equation that involved the total 940 tickets sold?

OpenStudy (theyankee):

Here's the explanation I was given. A : sold in adv D : sold day of concert A+D = 940 18A + 25D = 21820 (2) ==> D = 940 - A replace in (2) : 18A + 25D = 21820 ==> 18A + 25(940 - A) = 21820 ==> (25-18)A = 25*940 - 21820 ==> 7A = 1680 ==> A = 240 ==> D = 940 - 240 = 700 finally : A = 240 sold in advance D = 700 sold day of concert

OpenStudy (anonymous):

Ahhh okay. Perfect! That makes a lot more sense now. I really appreciate the help!

OpenStudy (theyankee):

Anytime :) I know how frustrating it can be to just not quite understand something. And I'll pass it onto him ^^

OpenStudy (anonymous):

Math man, I tell ya haha. I can write and learn history like no other, but for some reason my brain has trouble wrapping around the concepts of equations haha. Yes, please tell him thank you as well! =D

OpenStudy (theyankee):

I'm personally better with linear equations and in inequalities, but with effort, I can get these too. Tag me if you need anything else ^^